OpenStudy (anonymous):

Use Euler's formula to find the missing number. Vertices: 13 Edges: 28 Faces: ? 17 16 18 20 I'm so confused on this!!

OpenStudy (anonymous):

What is Euler's formula?

OpenStudy (anonymous):

F+V-E=2

OpenStudy (anonymous):

\[17+13=30,30-28=2\]

OpenStudy (anonymous):

Oh okay, I was mixed up because I was finding faces not vertices

OpenStudy (anonymous):

Oh, ok

OpenStudy (anonymous):

Can you help with more or?

OpenStudy (anonymous):

Are they more Euler's things, or something else?